# Continuum Transformers Perform In-Context Learning by Operator Gradient Descent

### v1

@@ 4 Experiments
We now wish to empirically verify the theoretically demonstrated claims. We summarize the relevant details of the setup in Section 4.1 and subsequently study the claims of Section 3.3 in Section 4.2 and those of Section 3.4 in Section 4.3. We additionally discuss E.4 for the chosen kernels in Section 4.3. Code for this verification will be made public upon acceptance.
@@ 4.1 Experimental Setup
In the experiments that follow, we wish to draw $\kappa$ Gaussian Random Variables, as per Definition 3.2. To begin, we first define an operator-valued covariance kernel $\kappa$. In particular, we consider one commonly encountered in the Bayesian functional data analysis literature [44], namely the Hilbert-Schmidt Integral Operator, defined as follows. Suppose $k_{x}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $k_{y}:\Omega\times\Omega\rightarrow\mathbb{R}$ are both positive definite kernels, where $k_{y}$ is Hermitian. Then, the following is a well-defined kernel:
@@
After fixing a choice of kernel, we then require a sampling procedure for operators from the RKHS induced by this choice of kernel. Notably, similar to how functions in a scalar RKHS can be sampled as $f=\sum_{s=1}^{S}\alpha^{(s)}k(x^{(s)},\cdot)$ for $\alpha^{(s)}\sim\mathcal{N}(0,\sigma^{2})$ over a randomly sampled collection $\{x_{i}\}_{i=1}^{n}$, we can sample operators by sampling a collection $\{(\varphi^{(s)},\psi^{(s)})\}_{s=1}^{S}$ and defining
@@
We focus on $\mathcal{X}=\mathcal{L}^{2}(\mathbb{T}^{2})$, for which the functions $\varphi^{(s)},\psi^{(s)}$ can be sampled from a Gaussian with mean function 0 and covariance operator $\alpha(-\Delta+\beta I)^{-\gamma}$, where $\alpha,\beta,\gamma\in\mathbb{R}$ are parameters that control the smoothness of the sampled functions and $\Delta$ is the Laplacian operator. Such a distribution is typical of the neural operators literature, as seen in [45, 46, 47], and is sampled as
$\varphi^{(s)}:=\sum_{|\nu|_{\infty}\leq N/2}\left(Z^{(s)}_{\nu}\alpha^{1/2}(4\pi^{2}||\nu||^{2}_{2}+\beta)^{-\gamma/2}\right)e^{i\nu\cdot x}\quad\text{ where }Z^{(s)}_{\nu}\sim\mathcal{N}(0,1),$ (17)
where $N/2$ is the Nyquist frequency assuming a discretized spatial resolution of $N\times N$. Such sampling is similarly repeated for $\psi^{(s)}$. We consider standard scalar kernels $k_{x}$ and $k_{y}$ to define Hilbert-Schmidt operator kernels. In particular, we consider the following kernel choices:
Kernels for $k_{x}(f,f^{\prime})$
Name Definition
Linear $\langle f,f^{\prime}\rangle_{\mathcal{X}}$
Kernels for $k_{y}(x,y)$
Name Definition
Laplace $\exp\left(-\frac{\|x-y\|_{2}}{\sigma_{y}}\right)$
Gaussian $\exp\left(-\frac{\|x-y\|_{2}^{2}}{2\sigma_{y}^{2}}\right)$
To finally construct the in-context windows, we similarly sample functions $f^{(j)}$ per Equation 17 and evaluate $u^{(j)}=Of^{(j)}$ using the sampled operator. To summarize, a single ICL context window $i$ of the form Equation 7 is constructed by defining an operator $O_{i}$ with a random sample $\{(\alpha_{i}^{(s)},\varphi_{i}^{(s)},\psi_{i}^{(s)})\}_{s=1}^{S}$, sampling input functions $\{f_{i}^{(j)}\}$, and evaluating $u_{i}^{(j)}=O_{i}f_{i}^{(j)}$.
4.2 Best Linear Unbiased Predictor Experiments
We now empirically demonstrate the claim of Proposition 3.3, specifically using the Hilbert-Schmidt operator-valued kernels described in the previous section. In particular, we demonstrate the optimality of the continuum transformer in-context update steps if the nonlinearity is made to match the kernel of the data-generating process. We consider four pairs of the aforementioned $(k_{x},k_{y})$ to define the operator-valued kernels and then fix the model parameters to be those given by Theorem 3.1. Notably, as $\mathcal{W}_{k,\ell}$ and $\mathcal{W}_{q,\ell}$ are implemented per Equation 2, we do so by fixing $R_{k,\ell}=R_{q,\ell}=\mathbbm{1}_{(N/2)\times(N/2)}$ $\forall\ell$. The results are shown in Figure 1. As expected, we find the in-context loss curves to decrease over increasing layers when the kernel matches the nonlinearity, as each additional layer then corresponds to an additional step of operator gradient descent per Theorem 3.1. We additionally see the optimality of matching the kernel and nonlinearity across the different choices of kernels in the limit of $\ell\rightarrow\infty$. We see that this result is robust over samplings of the operator, as the results in Figure 1 are reported over 50 independent trials. We additionally visualize the in-context learned predictions for each setup combination in Appendix F. As expected, we find that, when $k_{x}$ matches the data-generating kernel, the resulting field predictions structurally match the true $\mathcal{O}f$.
Figure 1: In-context learning loss curves over the number of layers in the continuum transformer. The kernels of the data-generating processes are given in the titles of the sub-figures. Curves show the mean ±$\pm 1/2$ standard deviation from 50 i.i.d. draws of the operator.
4.3 Optimization Experiments
Figure 2: Pairwise convergence of the key-key, key-query, and query-query operators of the continuum transformer in Hilbert-Schmidt cosine similarity across different kernels $k_{x}$ over training steps. Curves show the mean ±$\pm 1$ standard deviation from 5 i.i.d. trials of training procedure.
We now seek to empirically validate Theorem 3.6, namely that $\mathcal{W}_{k,\ell}=b_{\ell}\Sigma^{-1/2}$ and $\mathcal{W}_{q,\ell}=c_{\ell}\Sigma^{-1/2}$ are fixed points of the optimization landscape. Direct verification of this statement, however, is not feasible, since do not have an explicit form of the $\Sigma^{-1/2}$ operator. Nonetheless, we can verify that, for any $\ell_{1},\ell_{2}$ and $i,j\in\{q,k\}$ (indicating whether we are comparing a key-key, query-query, or key-query operator pair), $\langle\overline{\mathcal{W}}_{i,\ell_{1}},\overline{\mathcal{W}}_{j,\ell_{2}}\rangle_{\mathrm{HS}}\to 1$, where $\overline{\mathcal{W}}:=\mathcal{W}/||\mathcal{W}||_{\mathrm{HS}}$ denotes the normalized operator. Since we are working over the space $\mathcal{X}=\mathcal{L}^{2}(\mathbb{T}^{2})$ and considering Fourier kernel parameterization for the operators, this Hilbert-Schmidt inner product can be computed over the kernels, i.e. $\langle\overline{R}_{i,\ell_{1}},\overline{R}_{j,\ell_{2}}\rangle_{\mathrm{F}}$, where $\langle\cdot,\cdot\rangle_{\mathrm{F}}$ is the Frobenius inner product. The final metric is then
$\overline{\cos\theta}(t):=\frac{1}{4m^{2}}\sum_{i,j\in\{k,q\}}\sum_{\ell_{1},\ell_{2}\in\{1,...,m\}}\frac{\langle R_{i,\ell_{1}},R_{j,\ell_{2}}\rangle_{\mathrm{F}}}{||R^{(t)}_{i,\ell_{1}}||_{F}||R^{(t)}_{j,\ell_{2}}||_{F}},$ (18)
@@ 4.3 Optimization Experiments
As required by Theorem 3.6, we constrain the optimization of $\mathcal{W}_{v,\ell}$ to be over operators of the form $\begin{bmatrix}0&0\\ 0&r_{\ell}\end{bmatrix}$. This procedure is repeated across each of the $k_{x}$ kernels considered in the previous section with $k_{y}$ fixed to be the Gaussian kernel; we demonstrate that the Linear $k_{x}$ kernel satisfies E.4 in Section G.1. Notably, the other choices of $k_{x}$ do not satisfy this assumption, yet we find that the convergence result holds robustly to this violation. Specific hyperparameter choices of the training are presented in Appendix G. The results are presented in Figure 2. As mentioned, we see that the operators all converge pairwise on average, validating the characterization of the fixed points given by Theorem 3.6. As in the previous experiment, we find results to consistently replicate across training runs, pointing to this finding being robust to initializations.
@@ 5 Discussion
In this paper, we provided a theoretical characterization of the in-context learning phenomenon exhibited by continuum transformers and further validated such claims empirically. This work, therefore, naturally lends itself to several follow-up directions of interest. In the first, we would want to leverage our result to optimize in-context learning for PDEs, specifically by using the distribution induced on solution operators to inform transformer parameter choices per Theorem 3.1. Such operator distributions are often not explicitly defined but rather induced by distributions on parameters of the PDE and its parametric form. Additionally, the results of Section 4.3 suggest that a stronger version of Theorem 3.6 should hold, in which the convergence result is independent of $\ell$; proving such a generalization would be of great interest.
Appendix A RKHS Functional Gradient Descent Theorems
@@
Appendix B Gradient Descent in Operator Space
@@
B.1 Gradient Descent in Operator RKHS
@@
Definition B.2.
[…]
@@
    $\langle O\kappa(f,\cdot)u\rangle_{\mathcal{O}}=\langle Of,u\rangle_{\mathcal{Y}}$
@@
We now state the Representer Theorem for operator RKHS’s, as stated in Theorem 11 of [38]. Assume that $\mathcal{O}$ can be decomposed orthogonally into $\mathcal{O}=\mathcal{O}_{0}\oplus\mathcal{O}_{1}$ where $\mathcal{O}_{0}$ is a finite-dimensional Hilbert space spanned by the operators $\{E_{k}\}^{r}_{k=1}$ and $\mathcal{O}_{1}$ is its orthogonal complement under the inner product $\langle\cdot,\cdot\rangle_{\mathcal{O}}$. We denote the inner product $\langle\cdot,\cdot\rangle_{\mathcal{O}}$ restricted to $\mathcal{O}_{0},\mathcal{O}_{1}$ as $\langle\cdot,\cdot\rangle_{\mathcal{O}_{0}},\langle\cdot,\cdot\rangle_{\mathcal{O}_{0}}$ respectively.
Theorem B.3.
@@
$\widehat{O}(x,y)=\sum_{k=1}^{r}d_{k}E_{k}+\sum_{i=1}^{n}\alpha_{i}\kappa(f^{(i)},\cdot)$
for some $d\in\mathbb{R}^{r}$ and $\alpha_{i}\in\mathcal{Y}$.
@@
$G^{*}(\cdot)=c\sum_{i=1}^{n}\kappa(f^{(i)},\cdot)(u^{(i)}-Of^{(i)})$
@@
We apply theorem B.3 to equation Equation 27 with $\mathcal{O}_{0}$ the trivial subspace and $\psi(s)=\frac{\lambda}{2}s^{2}$. Then our solution has the form
@@
$\alpha^{*}=\frac{1}{\sum_{i,j=1}^{n}\langle\alpha_{i},\kappa(f^{(i)},f^{(j)})\alpha_{j}\rangle_{\mathcal{Y}}}(U-OF).$
@@
$G^{*}(\cdot)=\frac{1}{\sum_{i,j=1}^{n}\langle\alpha_{i},\kappa(f^{(i)},f^{(j)})\alpha_{j}\rangle_{\mathcal{Y}}}\sum_{i=1}^{n}\kappa(f^{(i)},\cdot)(u^{(i)}-Of^{(i)}).$
This gives us an exact form of $c$ as stated in equation Equation 29. We can re-write $c$ as
$c=\sum_{i,j=1}^{n}\langle\alpha_{i},\kappa(f^{(i)},f^{(j)})\alpha_{j}\rangle_{\mathcal{Y}}=\mathbf{1}^{T}M\mathbf{1}$
where $(M)_{ij}=\langle\alpha_{i},\kappa(f^{(i)},f^{(j)})\alpha_{j}\rangle_{\mathcal{Y}}$ and $\mathbf{1}$ is the vector of $1$’s. Then $c>0$ by the positive definite property of the kernel $\kappa$. ∎
Appendix C In-Context Learning via Gradient Descent Proof

### v2

@@ 5 Experiments
We now wish to empirically verify the theoretically demonstrated claims. We provide setup details in Section 5.1 and then study the claims of Section 3.3 in Section 5.2 and those of Section 3.4 in Section 5.3. Code for this verification will be made public upon acceptance.
@@ 5.1 Experimental Setup
In the experiments that follow, we wish to draw $\kappa$ Gaussian Random Variables, as per Definition 3.2. To begin, we first define an operator-valued covariance kernel $\kappa$. In particular, we consider one commonly encountered in the Bayesian functional data analysis literature [40], namely the Hilbert-Schmidt Integral Operator. Suppose $k_{x}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $k_{y}:\Omega\times\Omega\rightarrow\mathbb{R}$ are both positive definite kernels, where $k_{y}$ is Hermitian. Then, the following is a well-defined kernel:
$\kappa(f^{(1)},f^{(2)})u](y):=k_{x}(f^{(1)},f^{(2)})\int k_{y}(y^{\prime},y)u(y^{\prime})dy^{\prime}.$ (13)
Notably, similar to how functions in a scalar RKHS can be sampled as $f=\sum_{s=1}^{S}\alpha^{(s)}k(x^{(s)},\cdot)$ for $\alpha^{(s)}\sim\mathcal{N}(0,\sigma^{2})$ over a randomly sampled collection $\{x^{(s)}\}_{s=1}^{S}$, we can sample operators by sampling a collection $\{(\varphi^{(s)},\psi^{(s)})\}_{s=1}^{S}$ and defining $O=\sum_{i=s}^{S}\alpha^{(s)}\left(k_{x}(\varphi^{(s)},\cdot)\int k_{y}(y^{\prime},y)\psi^{(s)}(y^{\prime})dy^{\prime}\right)$ with $\alpha^{(s)}\sim\mathcal{N}(0,\sigma^{2})$. We focus on $\mathcal{X}=\mathcal{L}^{2}(\mathbb{T}^{2})$, for which the functions $\varphi^{(s)},\psi^{(s)}$ can be sampled from a Gaussian with mean function 0 and covariance operator $\alpha(-\Delta+\beta I)^{-\gamma}$, where $\alpha,\beta,\gamma\in\mathbb{R}$ are parameters that control the smoothness of the sampled functions and $\Delta$ is the Laplacian operator. Such a distribution is typical of the neural operators literature, as seen in [41, 42, 43], and is sampled as
@@
where $N/2$ is the Nyquist frequency assuming a discretized spatial resolution of $N\times N$. Such sampling is similarly repeated for $\psi^{(s)}$. We consider standard scalar kernels $k_{x}$ and $k_{y}$ to define Hilbert-Schmidt kernels. For $k_{x}$, we consider the Linear, Laplacian, Gradient RBF, and Energy kernels and for $k_{y}$, the Laplace and Gaussian kernels. The definitions of such kernels is deferred to Appendix H. To finally construct the in-context windows, we similarly sample functions $f^{(j)}$ per Equation 14 and evaluate $u^{(j)}=Of^{(j)}$ using the sampled operator. To summarize, a single ICL context window $i$ of the form Equation 6 is constructed by defining an operator $O_{i}$ with a random sample $\{(\alpha_{i}^{(s)},\varphi_{i}^{(s)},\psi_{i}^{(s)})\}_{s=1}^{S}$, sampling input functions $\{f_{i}^{(j)}\}$, and evaluating $u_{i}^{(j)}=O_{i}f_{i}^{(j)}$.
5.2 Best Linear Unbiased Predictor Experiments
Figure 1: In-context learning loss curves over the number of layers in the continuum transformer. The kernels of the data-generating processes are given in the titles of the sub-figures. Curves show the mean $\pm 1/2$ standard deviation from 50 i.i.d. draws of the operator.
We now empirically demonstrate the claim of Proposition 3.3, specifically using the Hilbert-Schmidt operator-valued kernels described in the previous section. In particular, we demonstrate the optimality of the continuum transformer in-context update steps if the nonlinearity is made to match the kernel of the data-generating process. We consider four pairs of the aforementioned $(k_{x},k_{y})$ to define the operator-valued kernels and then fix the model parameters to be those given by Theorem 3.1. Notably, as $\mathcal{W}_{k,\ell}$ and $\mathcal{W}_{q,\ell}$ are implemented as FNO kernel integral layers, we do so by fixing $R_{k,\ell}=R_{q,\ell}=\mathbbm{1}_{(N/2)\times(N/2)}$ $\forall\ell$. The results are shown in Figure 1. As expected, we find the in-context loss curves to decrease over increasing layers when the kernel matches the nonlinearity, as each additional layer then corresponds to an additional step of operator gradient descent per Theorem 3.1. We additionally see the optimality of matching the kernel and nonlinearity across the different choices of kernels in the limit of $\ell\rightarrow\infty$. We see that this result is robust over samplings of the operator, as the results in Figure 1 are reported over 50 independent trials. We visualize the in-context learned predictions for each setup combination in Appendix I, which reveals that, when $k_{x}$ matches the data-generating kernel, the resulting field predictions structurally match the true $\mathcal{O}f$.
5.3 Optimization Experiments
Figure 2: Pairwise convergence of the key-key, key-query, and query-query operators of the continuum transformer in Hilbert-Schmidt cosine similarity across different kernels $k_{x}$ over training steps. Curves show the mean $\pm 1$ standard deviation from 5 i.i.d. trials of training procedure.
@@ 5.3 Optimization Experiments
[…]
@@
As required by Theorem 3.6, we constrain the optimization of $\mathcal{W}_{v,\ell}$ to be over operators of the form $\begin{bmatrix}0&0\\ 0&r_{\ell}\end{bmatrix}$. This procedure is repeated across each of the $k_{x}$ kernels considered in the previous section with $k_{y}$ fixed to be the Gaussian kernel; we demonstrate that the Linear $k_{x}$ kernel satisfies G.4 in Section J.1. Notably, the other choices of $k_{x}$ do not satisfy this assumption, yet we find that the convergence result holds robustly to this violation. Specific hyperparameter choices of the training are presented in Appendix J. The results are presented in Figure 2. As mentioned, we see that the operators all converge pairwise on average, validating the characterization of the fixed points given by Theorem 3.6. As in the previous experiment, we find results to consistently replicate across training runs, pointing to this finding being robust to initializations.
@@ 6 Discussion
In this paper, we provided a theoretical characterization of the ICL phenomenon exhibited by continuum transformers and further validated such claims empirically. Unlike in the language learning context, this insight suggests a practical direction for improving ICL for meta-learning in PDEs, namely by estimating $\kappa$ for specific PDE meta-learning tasks and using this directly to parameterize $\widetilde{H}$ in the continuum transformer architecture. Such RKHSs are often not explicitly defined but rather induced by distributions on parameters of the PDE and its parametric form. Additionally, the results of Section 5.3 suggest that a stronger version of Theorem 3.6 should hold, in which the convergence result is independent of $\ell$; proving such a generalization would be of great interest.
Appendix A Continuum Transformer Token Details
In [12], the setting of interest was in extending the standard vision transformers that typically act on finite-dimensional images $\mathbb{R}^{H\times W}$ to instead act on a continuum, i.e. over a function mapping $\Omega\to\mathbb{R}$ for $\Omega\subset\mathbb{R}^{2}$. In doing so, they had to generalize the typical notions of patching that are introduced in vision transformers, by considering patches that decompose the domain, i.e. $\Omega_{i}\\$ } such that $\cup_{i}\Omega_{i}=\Omega$ and $\Omega_{i}\cap\Omega_{j}=\emptyset$. From here, each patch becomes a separate function $f_{i}:\Omega_{i}\to\mathbb{R}$, which necessitated dealing with infinite-dimensional tokens in their architecture, as we assumed directly for our setting.
Appendix B Time Rollout Meta Learning
As discussed in the main text, the pairs $(f^{(j)},u^{(j)})$ are generally time rollout pairs $(f^{(j)},u^{(j)})=(u^{((j-1)\Delta t)},u^{(j\Delta t)})$. We elaborate on this common setting below. In the context of doing time-rollouts, a pre-training dataset of the form $\mathcal{D}:=\{(U_{i}^{(0:T-1)},U_{i}^{(T)})\}$ is available, where $U_{i}^{(0:T)}:=[u_{i}^{(0)},u_{i}^{(\Delta T)},...,u_{i}^{((T-1)\Delta T)}]$, with $u_{i}^{(t)}\in\mathcal{L}^{2}(\Omega)$. Notably, such a setup is equivalent to having $n=T/\Delta T$ training pairs $\{(u_{i}^{((t-1)\Delta T)},u_{i}^{t\Delta T})\}_{t=1}^{n}$. It is further assumed that, for each sample $i$, there is a true, deterministic solution operator $\mathcal{G}_{i}\in\mathcal{O}$, where $\mathcal{O}$ is a space of operators, that maps from the spatial field at some time $t$ to its state at some later $t+\Delta T$. The in-context learning goal, thus, is, given a new sequence of $\widetilde{U}^{(0:T-1)}$ generated by some unseen operator $\widetilde{\mathcal{G}}$, predict $\widetilde{U}^{(T)}$, i.e.
$Z_{0}=\begin{pmatrix}\tilde{u}^{(0)}&\tilde{u}^{(\Delta t)}&\ldots&\tilde{u}^{((n-1)\Delta t)}&\tilde{u}^{(n\Delta t)}\\ \\ \tilde{u}^{(\Delta t)}&\tilde{u}^{(2\Delta t)}&\ldots&\tilde{u}^{(n\Delta t)}&0\end{pmatrix}$
Appendix C RKHS Functional Gradient Descent Theorems
@@
Appendix D Gradient Descent in Operator Space
@@
D.1 Gradient Descent in Operator RKHS
@@
Definition D.2.
Operator RKHS Let $\mathcal{O}$ be a Hilbert space of operators $O:\mathcal{X}\to\mathcal{Y}$, equipped with an inner product $\langle\cdot,\cdot\rangle_{\mathcal{O}}$. We call $\mathcal{O}$ an operator RKHS if there exists an operator-valued kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{L}(\mathcal{Y})$ such that
@@
    $\langle O,\kappa(f,\cdot)u\rangle_{\mathcal{O}}=\langle Of,u\rangle_{\mathcal{Y}}$
@@
We now state the Representer Theorem for operator RKHS’s, as stated in Theorem 11 of [45]. Assume that $\mathcal{O}$ can be decomposed orthogonally into $\mathcal{O}=\mathcal{O}_{0}\oplus\mathcal{O}_{1}$ where $\mathcal{O}_{0}$ is a finite-dimensional Hilbert space spanned by the operators $\{E_{k}\}^{r}_{k=1}$ and $\mathcal{O}_{1}$ is its orthogonal complement under the inner product $\langle\cdot,\cdot\rangle_{\mathcal{O}}$. We denote the inner product $\langle\cdot,\cdot\rangle_{\mathcal{O}}$ restricted to $\mathcal{O}_{0},\mathcal{O}_{1}$ as $\langle\cdot,\cdot\rangle_{\mathcal{O}_{0}},\langle\cdot,\cdot\rangle_{\mathcal{O}_{1}}$ respectively.
Theorem D.3.
@@
$\widehat{O}=\sum_{k=1}^{r}d_{k}E_{k}+\sum_{i=1}^{n}\kappa(f^{(i)},\cdot)\alpha_{i}$
for some $d\in\mathbb{R}^{r}$, $E_{k}\in\mathcal{O}_{0}$ for all $k\in\{1,\ldots,r\}$ and $\alpha_{i}\in\mathcal{Y}$ for all $i\in\{1,\ldots,n\}$.
@@
$G^{*}(\cdot)=c\sum_{i=1}^{n}\kappa(f^{(i)},\cdot)(u^{(i)}-Of^{(i)})$ (26)
@@
We apply theorem D.3 to equation Equation 24 with $\mathcal{O}_{0}$ the trivial subspace and $\psi(s)=\frac{\lambda}{2}s^{2}$. Then our solution has the form
@@
$\alpha^{*}=\frac{1}{\sum_{i,j=1}^{n}\langle u^{(i)}-Of^{(i)},\kappa(f^{(i)},f^{(j)})(u^{(j)}-Of^{(j)})\rangle_{\mathcal{Y}}}(U-OF).$
@@
$G^{*}(\cdot)=\frac{1}{\sum_{i,j=1}^{n}\langle u^{(i)}-Of^{(i)},\kappa(f^{(i)},f^{(j)})(u^{(j)}-Of^{(j)})\rangle_{\mathcal{Y}}}\sum_{i=1}^{n}\kappa(f^{(i)},\cdot)(u^{(i)}-Of^{(i)}).$
This gives us an exact form of $c$ as stated in equation Equation 26. ∎
Appendix E In-Context Learning via Gradient Descent Proof